### v1

# Experiment design for continuous-time systems using generalized filtering

Abstract
The goal of experiment design is to select the inputs of a dynamical system in such a way that the resulting data contain sufficient information for data-driven modeling and control. This paper investigates the problem of experiment design for continuous-time systems under piecewise constant input signals. To obviate the need for measuring time derivatives of (data) trajectories, we introduce a generalized filtering framework. Our main result is to establish conditions on the input and the filter functions under which the filtered data are informative for system identification, i.e., they satisfy a certain rank condition. The proposed framework encompasses several filter functions that have already appeared in the literature. Building on the proposed filtering framework, we develop an experiment design procedure where the input signal is designed online during system operation. This method is shown to be sample efficient, in the sense that it deals with the least possible number of filtered data samples for system identification.
keywords:
Continuous-time systems, experiment design, fundamental lemma, generalized filtering, system identification.
thanks: This paper was not presented at any IFAC meeting. Corresponding author: Simone Baldi. This work was partially supported by Jiangsu Provincial Scientific Research Center of Applied Mathematics grant BK20233002. H. J. van Waarde acknowledges financial support by the Dutch Research Council under the NWO Talent Programme Veni Agreement (VI.Veni.222.335).
1 Introduction
The growing complexity of modern engineering systems complicates modeling from first principles and motivates the use of data for system modeling and control [de2019formulas, van2023informativity, DBLSCT2025]. Data-driven techniques utilize measurements collected from the system to model the system itself or to design controllers. Data-driven techniques for discrete-time systems [de2019formulas, van2023informativity, DBLSCT2025, van2020data, berberich2021data, rotulo2022online, wang2025necessary, eising2025data, zhao2025data] have a richer development as compared to their continuous-time counterparts [berberich2021datacontinuous, rapisarda2023orthogonal, ohta2024data, eising2025sampling], partly due to the challenge of accurately measuring time derivatives of trajectories in continuous-time systems. One of the possible approaches to address this challenge is to apply one of the various filtering methods proposed in the literature [ohsumi2002subspace, ohta2011stochastic, ohta2024data]. Alternative methods involve, for instance, robust adaptation [yucelen2011derivative], estimation using discretization [berberich2021datacontinuous] and orthogonal polynomial bases [rapisarda2023orthogonal].
While most of the existing literature performs analysis and control using given datasets, other studies address a preliminary question—the design of input signals, with the aim of generating suitable data. This phase, known as experiment design, is crucial to obtain data containing sufficient information for data-driven modeling and control. The theoretical foundations for experiment design are provided by the fundamental lemma by Willems et al., originally formulated for discrete-time systems [willems2005note]. The lemma characterizes those input signals resulting in collected data that represent all possible trajectories of the system. Different extensions of this lemma to continuous-time systems have been proposed [rapisarda2022persistency, lopez2022continuous, rapisarda2023fundamental, lopez2024input], demonstrating from different perspectives that a single persistently exciting input signal can generate a system trajectory rich enough to reproduce all other trajectories. Nevertheless, these results typically rely on direct access to a number of time derivatives of (data) trajectories. The experiment design for continuous-time systems without the availability of time derivatives remains a challenge.
This naturally raises the question of whether such a problem can be addressed by leveraging on filtering methods like those in [ohsumi2002subspace, ohta2011stochastic, ohta2024data]. This work gives a positive answer to this question by means of a generalized filtering framework encompassing existing filtering methods. Answering this question is far from trivial, as no existing result in the literature guarantees that the data obtained after filtering will retain sufficient information for data-driven modeling and control. In this work, we provide conditions under which the filtered data are informative for system identification, i.e., they satisfy a certain rank condition.
Throughout this paper, we consider two types of data: sampled data, obtained from the continuous-time system through piecewise constant input signals and sampling of the trajectories, and filtered data, obtained by filtering the trajectories via generalized filter functions. Based on this, we address two key problems:
  • 1)
    Find conditions on the trajectories and the filter functions under which the set of filtered data is informative for system identification.
  • 2)
    Develop an experiment design method that generates piecewise constant inputs and guarantees that the resulting filtered data are informative.
In this work, we introduce a generalized filtering framework (cf. Eq. (1)) which encompasses several filtering methods from the literature, see Example 1. Based on this framework, the two main contributions of the paper are as follows:
  • We establish a relation between the filtered data and the sampled data (cf. Eq. (32)). This relation enables us to establish Proposition 1, which provides conditions under which the filtered data are informative for system identification. Notably, the conditions we present are missing in all filtering methods in the literature, which rely on assuming informativity of the filtered data, rather than deriving the conditions making informativity possible.
  • We propose an online experiment design method (cf. Proposition 3) that guarantees the filtered data to be informative with the minimum possible number of samples. Theorem 2 shows that the designed input signal is such that the data captures the system’s dynamics at all times between sampling instants, which establishes a connection between the proposed method and the continuous-time Willems et al.’s fundamental lemma in [lopez2022continuous].
The remainder of this paper is organized as follows. Section 2 introduces the generalized filtering framework. The problems addressed in the paper are formulated in Section 3. Section 4 investigates the relation between the filtered data and the sampled data. The proposed online experiment design method is discussed in Section 5. Section 6 provides a numerical illustration of the theoretical development. Finally, Section 7 concludes the paper.
1.1 Notation
We denote the set of non-negative integers by $\mathbb{Z}_{+}$, the set of positive integers by $\mathbb{N}$, and the set of non-negative real numbers by $\mathbb{R}_{+}$. Given a matrix $A\in\mathbb{R}^{n\times m}$, its Moore-Penrose pseudo-inverse is denoted by $A^{\dagger}$. The identity matrix with appropriate dimensions is denoted by $I$. Given vectors $v_{k},v_{k+1},\dots,v_{\ell}\in\mathbb{R}^{n}$ with $k\leq\ell$, we define $v_{[k,\ell]}:=[v_{k}\ v_{k+1}\ \cdots\ v_{\ell}]\in\mathbb{R}^{n\times(\ell-k+1)}$.
2 Generalized filtering and filtered data
For $\mathcal{T}>0$ and $\ell\in\{1,2,\dots,M\}$, we introduce the filter functions $g_{\ell}:[0,\mathcal{T})\to\mathbb{R}$. We let $0=t_{0}<t_{1}<t_{2}<\cdots<t_{q}=\mathcal{T}$ and assume that $g_{\ell}$ is continuously differentiable on $t_{j-1},t_{j})$ for $j=1,2,\dots,q$. Thus, $g_{\ell}$ is a piecewise continuously differentiable function. In addition, we will assume that $g_{\ell}(t_{j}^{-}):=\lim_{t\uparrow t_{j}}g_{\ell}(t)$ exists for all $j=1,2,\dots,q$. The functions $g_{\ell}$ are used to produce filtered data $w^{\rm f}_{\ell}$ from a given signal $w:[0,\mathcal{T})\to\mathbb{R}^{n}$. More precisely, we define
$w^{\rm f}_{\ell}:=\int_{0}^{\mathcal{T}}g_{\ell}(t)w(t)dt,\quad\ell\in\{1,2,\dots,M\}.$ (1)
Note that (1) can accommodate several filtering methods from the literature as special cases. For instance, in [ohta2024data], the following set of filtered data was considered:
$\mathcal{D}=\{\langle S_{\tau}h,w\rangle\!:\!h\!\in\!\{h_{1},h_{2},\dots,h_{p}\},\tau\!\in\!\{s_{1},s_{2},\dots,s_{r}\}\},$
where
$\langle S_{\tau}h,w\rangle=\int_{\tau}^{\mathcal{T}}h(t-\tau)w(t)\,dt,$
and $h_{j}:[0,\mathcal{T})\to\mathbb{R}$ and $s_{k}\in[0,\mathcal{T})$ for $j=1,2,\dots,p$ and $k=1,2,\dots,r$. Note that $\mathcal{D}$ can be captured by (1) by setting $M=pq$ and defining
[…]
for $j\in\{1,\dots,p\}$ and $k\in\{1,\dots,q\}$. The generality of (1) is further elaborated in the next example.
Example 1.
(Filter functions). Several filter functions $g_{\ell}$ considered in the literature can be used in (1). For $T\geq 0$ and $\rho>0$, examples include:
  • i)
    compactly supported test functions [ohsumi2002subspace] such as:
    $g_{\ell}(t)\!=\!\left\{\begin{aligned} &\!\rho(t\!-\!(\ell\!-\!1)T)^{2}(\ell T\!-\!t)^{2}&&\!\!\text{if }t\!\in\![(\ell\!-\!1)T,\ell T),\\ &\!0&&\!\!\text{otherwise;}\\ \end{aligned}\right$. (2)
  • or
[…]
  • ii)
    first Laguerre basis function [ohta2024data]:
[…]
  • iii)
    low-pass filter function [roy2017combined]:
[…]
(a) $g_{\ell}(t)=(t-\ell+1)^{2}(\ell-t)^{2}$
(b) $g_{\ell}(t)=e^{-\frac{1}{1-(t-\ell+1)^{2}}}$
(c) $g_{\ell}(t)=\sqrt{2}e^{-(t-\ell+1)}$
(d) $g_{\ell}(t)=e^{t-\ell}$
Figure 1: Examples of filter functions for $\ell\in\{1,2,3\}$
Fig. 1(a)-1(d) show the four filter functions in (2)-(5) with parameters $\rho=1$, $T=1$ and $\ell\in\{1,2,3\}$. In some cases, the filtered data in (1) can be interpreted as the state of a dynamical system sampled at certain time instants. To illustrate this, consider the low-pass filter function (5). For any signal $w:[0,\mathcal{T})\to\mathbb{R}^{n}$, low-pass filtering can be described by the differential equation
$\dot{w}^{\rm f}(t)=-\rho{w}^{\rm f}(t)+w(t),\quad w^{\rm f}(0)=0,$ (6)
where $w^{\rm f}(t)\in\mathbb{R}^{n}$ denotes the filtered signal. The solution of (6) is
$w^{\rm f}(t)=\int_{0}^{t}e^{-\rho(t-\tau)}w(\tau)d\tau.$ (7)
Hence, the following relation holds:
$w^{\rm f}(\ell T)=\int_{0}^{\ell T}\!e^{\rho(\tau-\ell T)}w(\tau)d\tau=\int_{0}^{\mathcal{T}}\!g_{\ell}(\tau)w(\tau)d\tau=w_{\ell}^{\rm f},$ (8)
where we have used (1), and the specific low pass filtering function $g_{\ell}$ defined in (5). We conclude that the filtered data $w_{\ell}^{f}$ can be interpreted as the state of the dynamical system (6) at specific time instants. The low-pass filter function (5) is often employed to avoid the need for measuring the time derivative of (state) variables [roy2017combined, cho2017composite]. Indeed, denote
$w^{\rm df}_{\ell}=\int_{0}^{\mathcal{T}}g_{\ell}(\tau)\dot{w}(\tau)d\tau\in\mathbb{R}^{n}.$
Then, using integration by parts, we obtain
$w^{\rm df}_{\ell}$ $=\int_{0}^{\ell T}e^{\rho(\tau-\ell T)}\dot{w}(\tau)d\tau$ (9)
$=w(\ell T)-e^{-\rho\ell T}w(0)-\rho w^{\rm f}_{\ell},$
i.e., the filtered time derivative data $w^{\rm df}_{\ell}$ can be calculated without measuring the time derivative $\dot{w}$.
3 Problem formulation
The filtering approach in (1) is now applied to signals generated by a continuous-time system. Consider the continuous-time linear time-invariant system
$\dot{x}(t)$ $=Ax(t)+Bu(t),$ (10a)
$x(0)$ $=x_{0},$ (10b)
where $x(t)\in\mathbb{R}^{n}$ is the system state, $u(t)\in\mathbb{R}^{m}$ is the control input, and $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times m}$ are the system matrices, assumed to be unknown. Throughout the paper, we do assume that the pair $(A,B)$ is controllable.
Let $\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ be the space of absolutely continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{n}$, and let $\mathcal{PC}([0,\mathcal{T}),\mathbb{R}^{m})$ be the space of piecewise continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{m}$. We recall that $x\in\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ is a Carathéodory solution of (10) on $0,\mathcal{T})$ if (10a) holds for almost all $t\in[0,\mathcal{T})$ and (10b) is satisfied. We define the behavior [willems1997introduction] of (10) as
$\mathfrak{B}_{\mathcal{T}}:=\{$ $(x,u)\in\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})\times\mathcal{PC}([0,\mathcal{T}),\mathbb{R}^{m})\mid$
[…]
We call elements of $\mathfrak{B}_{\mathcal{T}}$ (input-state) trajectories of (10) on $0,\mathcal{T})$.
Given a trajectory $(u,x)\in\mathfrak{B}_{\mathcal{T}}$, we denote the collection of filtered data by
$x^{\rm df}_{[1,M]}$= $\begin{bmatrix}x^{\rm df}_{1}&x^{\rm df}_{2}&\cdots&x^{\rm df}_{M}\end{bmatrix}\in\mathbb{R}^{n\times M},$ (11)
$x^{\rm f}_{[1,M]}$= $\begin{bmatrix}x^{\rm f}_{1}&x^{\rm f}_{2}&\cdots&x^{\rm f}_{M}\end{bmatrix}\in\mathbb{R}^{n\times M},$
$u^{\rm f}_{[1,M]}$= $\begin{bmatrix}u^{\rm f}_{1}&u^{\rm f}_{2}&\cdots&u^{\rm f}_{M}\end{bmatrix}\in\mathbb{R}^{m\times M},$
where
$x^{\rm df}_{\ell}$ $=\int_{0}^{\mathcal{T}}g_{\ell}(\tau)\dot{x}(\tau)d\tau\in\mathbb{R}^{n},$ (12)
$x^{\rm f}_{\ell}$ $=\int_{0}^{\mathcal{T}}g_{\ell}(\tau)x(\tau)d\tau\in\mathbb{R}^{n},$ (13)
$u^{\rm f}_{\ell}$ $=\int_{0}^{\mathcal{T}}g_{\ell}(\tau)u(\tau)d\tau\in\mathbb{R}^{m},$ (14)
for $\ell\in\{1,2,\dots,M\}$. Then, (12) can be further rewritten using integration by parts:
$x^{\rm df}_{\ell}=\sum_{j=1}^{q}\int_{t_{j-1}}^{t_{j}^{-}}g_{\ell}(\tau)\dot{x}(\tau)d\tau=$ (15)
$\!\sum_{j=1}^{q}\!\Bigg(g_{\ell}(t_{j}^{-})x(t_{j}^{-})\!-\!g_{\ell}(t_{j-1})x(t_{j-1})\!-\!\int_{t_{j-1}}^{t_{j}^{-}}\!\dot{g}_{\ell}(\tau)x(\tau)d\tau\Bigg),$
where $x(t_{j}^{-})$ is well defined for all $j\in\{1,2,\dots,q\}$ because $x$ is absolutely continuous. This shows that there is no need for measuring the time derivative $\dot{x}$ of the state in order to obtain the filtered time derivative $x^{\rm df}_{\ell}$. Indeed, by (15), $x_{\ell}^{\rm df}$ may be obtained directly from measurements of $x$ and the filter function $g_{\ell}$. Since integration is a linear transformation, (12)-(14) and (10) yield the algebraic equation
$x^{\rm df}_{[1,M]}=Ax^{\rm f}_{[1,M]}+Bu^{\rm f}_{[1,M]}.$ (16)
In this work, we will consider the rank condition
$\operatorname{rank}\left(\begin{bmatrix}x^{\rm f}_{[1,M]}\\ u^{\rm f}_{[1,M]}\end{bmatrix}\right)=n+m.$ (17)
The rank condition (17) is crucial for data-driven modeling and control based on filtered data, see e.g., [ohta2024data]. Indeed, as an example we note that (16) and (17) imply that the system matrices $A$ and $B$ can be uniquely recovered from the data as
$\begin{bmatrix}A&B\end{bmatrix}=x^{\rm df}_{[1,M]}\begin{bmatrix}x^{\rm f}_{[1,M]}\\ u^{\rm f}_{[1,M]}\end{bmatrix}^{\dagger},$ (18)
i.e., the filtered data $(x^{\rm df}_{[1,M]},x^{\rm f}_{[1,M]},u^{\rm f}_{[1,M]})$ are informative for system identification. Analogous versions of the rank condition in (17) have also appeared for other types of data in [de2019formulas, berberich2021datacontinuous, rapisarda2023orthogonal], for example, for samples of continuous-time trajectories [de2019formulas, berberich2021datacontinuous] and coefficient matrices corresponding to certain basis functions [rapisarda2023orthogonal].
In this paper, our goal will be to establish conditions on $u$ such that the filtered data (11) resulting from the trajectory $(u,x)\in\mathfrak{B}_{\mathcal{T}}$ satisfy the rank condition (17). This leads to the formulation of the following two problems.
Problem 1.
Consider the trajectory $(u,x)\in\mathfrak{B}_{\mathcal{T}}$ and the filter functions $g_{\ell}$ for $\ell=1,2,\dots,M$. Provide conditions on this trajectory and filter functions such that (17) holds.
Problem 2.
Consider the filter functions $g_{\ell}$ for $\ell=1,2,\dots,M$. Design the input signal $u:[0,\mathcal{T})\to\mathbb{R}^{m}$ such that for every initial state $x(0)=x_{0}\in\mathbb{R}^{n}$, the resulting trajectory $(u,x)\in\mathfrak{B}_{\mathcal{T}}$ is such that the filtered data $(u^{f}_{[1,M]},x^{f}_{[1,M]})$ satisfy (17).
Note that for (17) to hold, it is necessary that $M\geq n+m$. In this paper, we are especially interested in the case that $M=n+m$ because it corresponds to the minimum number of filtered data samples required to achieve (17). It will follow from the results of this paper that there indeed exist inputs achieving (17) for $M=n+m$.
In the next two sections, we solve Problem 1 and Problem 2, respectively.
4 Rank condition for filtered data
Let $T>0$ and $N\in\mathbb{N}$ and define $\mathcal{T}:=NT$. We consider a piecewise constant input $u:[0,\mathcal{T})\to\mathbb{R}^{m}$, i.e.,
$u(t+kT)=\mu_{k},$ (19)
where $t\in[0,T)$ and $\mu_{k}\in\mathbb{R}^{m}$ for $k=0,1,\dots,N-1$. For the trajectory $(u,x)\in\mathfrak{B}_{NT}$, we consider sampled data at the time instants $t+kT$, captured in the matrices
$\chi_{[0,N-1]}(t):=\begin{bmatrix}\chi_{0}(t)&\chi_{1}(t)&\cdots&\chi_{N-1}(t)\end{bmatrix}\in\mathbb{R}^{n\times N},$ (20)
where $\chi_{k}(t)=x(t+kT)$. We use the shorthand notation $\chi_{[0,N-1]}:=\chi_{[0,N-1]}(0)$ and we denote
$\mu_{[0,N-1]}:=\begin{bmatrix}\mu_{0}&\mu_{1}&\cdots&\mu_{N-1}\end{bmatrix}\in\mathbb{R}^{m\times N}.$ (21)
4.1 Continuous-time fundamental lemma
For a piecewise constant input, exact discretization of the continuous-time system (10) with sampling time $T$ results in the discrete-time dynamics [chen2012optimal]
$\chi_{k+1}=A_{T}\chi_{k}+B_{T}\mu_{k},$ (22)
where
$A_{T}:=e^{AT},\quad B_{T}:=\int_{0}^{T}e^{At}Bdt.$ (23)
In order to ensure that the discrete-time system (22) preserves the controllability property of the original continuous-time system (10), the following assumption from the literature is considered.
Assumption 1.
Let $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$ be the eigenvalues of $A$. Then, for any distinct $j,l\in\{1,2,\dots,n\}$ and $q\in\mathbb{N}$, the sampling time $T$ is such that
$\lambda_{j}-\lambda_{l}\neq\frac{2q\pi}{T}i,$ (24)
where $i$ is the imaginary unit.
As highlighted in [chen2012optimal], Assumption 1 is mild in the sense that (24) holds for all $T$ except for a set of measure zero. Based on Assumption 1, we recall a lemma asserting that controllability is preserved under discretization.
Lemma 1.
([chen2012optimal, Theorem 3.2.1]). Let Assumption 1 hold. Then $(A_{T},B_{T})$ is controllable if $(A,B)$ is controllable.
Using this property, we recall the continuous-time fundamental lemma from [lopez2022continuous].
Lemma 2.
([lopez2022continuous, Lemma 1]). Let Assumption 1 hold. Consider the trajectory $(u,x)\in\mathfrak{B}_{NT}$ with $u$ a piecewise constant input signal in (19) satisfying $\operatorname{rank}(\mathcal{H}_{n+1}(\mu_{[0,N-1]}))=(n+1)m,$, where
$\mathcal{H}_{n+1}(\mu_{[0,N-1]})=\begin{bmatrix}\mu_{0}&\mu_{1}&\cdots&\mu_{N-n-1}\\ \vdots&\vdots&&\vdots\\ \mu_{n}&\mu_{n+1}&\cdots&\mu_{N-1}\\ \end{bmatrix}$,
is the Hankel matrix of $\mu_{[0,N-1]}$ of depth $n+1$. Then,
$\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,N-1]}(t)\\ \mu_{[0,N-1]}\end{bmatrix}\right)=n+m,\quad\forall t\in[0,T).$ (25)
Lemma 2 provides an experiment design method to achieve (25) for the sampled data (20)-(21). Yet, we stress on the fact that nothing can be concluded about the rank condition (17) for the filtered data, as required to solve Problem 1 and 2. Hence, we proceed by analyzing the relation between the sampled and the filtered data under piecewise constant input signals.
4.2 Rank relation between sampled and filtered data
We aim to establish a rank relation between the sampled and the filtered data. To this end, we first analyze the filtered state and filtered input data. By applying the piecewise constant input in (19), it follows from (13) and (14) that for any $\ell\in\{1,2,\dots,M\}$,
$x_{\ell}^{\rm f}=$ $\sum_{j=0}^{N-1}\int_{(j-1)T}^{jT}g_{\ell}(\tau)x(\tau)d\tau$
$=$ $\sum_{j=0}^{N-1}\int_{0}^{T}g_{\ell}(\tau+jT)x(\tau+jT)d\tau$
$=$ $\sum_{j=0}^{N-1}\int_{0}^{T}g_{\ell}(\tau+jT)e^{A\tau}d\tau\chi_{j}+$
$\sum_{j=0}^{N-1}\int_{0}^{T}g_{\ell}(\tau+jT)\int_{0}^{\tau}e^{A(\tau-s)}dsd\tau B\mu_{j},$
$u_{\ell}^{\rm f}=$ $\sum_{j=0}^{N-1}\int_{(j-1)T}^{jT}g_{\ell}(\tau)u(\tau)d\tau=\sum_{j=0}^{N-1}\int_{0}^{T}g_{\ell}(\tau+jT)d\tau\mu_{j}.$
Now, to build the relation between $x_{\ell}^{\rm f}$ and $\chi_{[0,N-1]}$, we make the following assumption.
Assumption 2.
For all $\tau\!\in\![0,T)$ and $j\!\in\!\{0,1,\dots,N\!-\!1\}$, the filter function $g_{\ell}$ satisfies
$g_{\ell}(\tau+jT)=g(\tau)f_{\ell}(jT),$ (26)
where $g:[0,T)\to[0,\infty)$ is a continuously differentiable function such that $\int_{0}^{T}g(\tau)d\tau>0$, and $f_{\ell}:\mathbb{R}\to\mathbb{R}$ for $\ell\in\{1,\dots,M\}$.
Example 2.
(Validity of Assumption 2). Suppose that $N\geq M$. A decomposition as in (26) can be found for all the filter functions in Example 1, namely,
  • i)
    compactly supported test functions:
    $g(t)=\rho t^{2}(T-t)^{2}\text{ or }g(t)=e^{-\frac{bT^{2}}{T^{2}-t^{2}}},$ (27)
[…]
  • ii)
    first Laguerre basis function:
    $g(t)=\sqrt{2\rho}e^{-\rho t},$ (28)
[…]
  • iii)
    low-pass filter function:
    $g(t)=e^{\rho t},$ (29)
[…]
For all cases, $g(t)\!\geq\!0$ for any $t\!\in\![0,T)$ and $\int_{0}^{T}\!g(\tau)d\tau\!>\!0$.
Under Assumption 2, we have
$x_{\ell}^{\rm f}=\sum_{j=0}^{N-1}\bar{A}\chi_{j}f_{\ell}(jT)+\sum_{j=0}^{N-1}\bar{B}\mu_{j}f_{\ell}(jT),$ (30)
$u_{\ell}^{\rm f}=\sum_{j=0}^{N-1}\bar{G}\mu_{j}f_{\ell}(jT),$ (31)
where
$\bar{A}$ $=\int_{0}^{T}g(\tau)e^{A\tau}d\tau\in\mathbb{R}^{n\times n},$
$\bar{B}$ $=\int_{0}^{T}g(\tau)\int_{0}^{\tau}e^{A(\tau-s)}Bdsd\tau\in\mathbb{R}^{n\times m},$
$\bar{G}=I\cdot\int_{0}^{T}g(\tau)d\tau\in\mathbb{R}^{m\times m}.\quad$
This results in the relation
$\begin{bmatrix}x_{[1,M]}^{f}\\ u_{[1,M]}\end{bmatrix}=\bar{C}\bar{D}\bar{F},$ (32)
where
$\bar{C}=$ $\begin{bmatrix}\bar{A}&\bar{B}\\ 0&\bar{G}\end{bmatrix},\quad\bar{D}=\begin{bmatrix}\chi_{[0,N-1]}\\ \mu_{[0,N-1]}\end{bmatrix}$,
$\bar{F}=$ $\begin{bmatrix}f_{1}(0)&f_{2}(0)&\cdots&f_{M}(0)\\ f_{1}(T)&f_{2}(T)&\cdots&f_{M}(T)\\ \vdots&\vdots&\ddots&\vdots\\ f_{1}((N-1)T)&f_{2}((N-1)T)&\cdots&f_{M}((N-1)T)\\ \end{bmatrix}\!$.
Now, to understand under which conditions the rank condition (17) holds, we first present two lemmas.
Lemma 3.
We have that
$\operatorname{rank}(\bar{C})=n+m.$ (33)
Proof. We first show that $\bar{A}$ is nonsingular. We consider the Jordan normal form of $A$, i.e., $A=P\Lambda P^{-1}$, where $P\in\mathbb{R}^{n\times n}$ is nonsingular,
$\Lambda=\!\begin{bmatrix}\Lambda_{1}&0&\cdots&0\\ 0&\Lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Lambda_{k}\end{bmatrix}\!,\ \Lambda_{l}=\!\begin{bmatrix}\lambda_{l}&1&\cdots&0&0\\ 0&\lambda_{l}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\lambda_{l}&1\\ 0&0&\cdots&0&\lambda_{l}\\ \end{bmatrix}\!\in\mathbb{R}^{n_{l}\times n_{l}},$
$l\in\{1,2,\dots,q\}$ and $\sum_{l=1}^{q}n_{l}=n$. Then,
$e^{At}=\sum_{j=0}^{\infty}\frac{A^{j}}{j!}t^{j}=P\sum_{j=0}^{\infty}\frac{\Lambda^{j}}{j!}t^{j}P^{-1}=Pe^{\Lambda t}P^{-1},$
implying that
$\det\left(\int_{0}^{T}g(\tau)e^{A\tau}d\tau\right)$ (34)
$=$ $\det(P)\det(\int_{0}^{T}g(\tau)e^{\Lambda\tau}d\tau)\det(P^{-1})$
$=$ $\prod_{{l}=1}^{q}\det\left(\int_{0}^{T}g(\tau)e^{\Lambda_{l}\tau}d\tau\right).$
Note that for any ${l}\in\{1,2,\dots,q\}$,
$g(\tau)e^{\Lambda_{l}\tau}=\begin{bmatrix}g(\tau)e^{\lambda_{l}\tau}&*&\cdots&*\\ 0&g(\tau)e^{\lambda_{l}\tau}&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&g(\tau)e^{\lambda_{l}\tau}\\ \end{bmatrix}$,
where $\ast$ denotes an entry that is left unspecified. By Assumption 2, there exist $t_{1},t_{2}\in(0,T)$ such that $t_{1}<t_{2}$ and $g(t)>0$ for any $t\in[t_{1},t_{2}]$. Then, since $g(\tau)e^{\lambda_{l}\tau}\geq 0$ for any $\tau\in[0,T)$, we have $\int_{0}^{T}g(\tau)e^{\lambda_{l}\tau}d\tau\geq\int_{t_{1}}^{t_{2}}g(\tau)e^{\lambda_{l}\tau}d\tau>0$, implying $\det(\int_{0}^{T}g(\tau)e^{\Lambda_{l}\tau}d\tau)\neq 0$. Therefore, by (34), we have that $\bar{A}$ is nonsingular. By Assumption 2, $\bar{G}$ is nonsingular, implying that $\bar{C}$ is nonsingular, i.e., (33) holds. $\Box$
Lemma 4.
([bernstein2009matrix, Fact 2.10.14]) For any $P\in\mathbb{R}^{p\times q}$ and $Q\in\mathbb{R}^{q\times r}$,
$\operatorname{rank}(PQ)=\operatorname{rank}(Q)-\dim(\ker(P)\cap\operatorname{im}(Q)).$ (35)
Now we investigate under which conditions (17) holds. By Lemma 3, $\bar{C}$ is nonsingular. Therefore, (32) implies
$\operatorname{rank}\left(\begin{bmatrix}x^{\rm f}_{[1,M]}\\ u^{\rm f}_{[1,M]}\end{bmatrix}\right)=\operatorname{rank}(\bar{D}\bar{F}),$
that is, (17) holds if and only if $\operatorname{rank}(\bar{D}\bar{F})=n+m$. By Lemma 4, we have
$\operatorname{rank}(\bar{D}\bar{F})=\operatorname{rank}(\bar{F})-\dim(\ker(\bar{D})\cap\operatorname{im}(\bar{F})).$
To arrive at the condition $\operatorname{rank}(\bar{D}\bar{F})=n+m$, we therefore require that
$\operatorname{rank}(\bar{F})-\dim(\ker(\bar{D})\cap\operatorname{im}(\bar{F}))=n+m.$
The latter condition is, in general, difficult to impose by choosing the inputs $\mu_{0},\mu_{1},....,\mu_{N-1}$. Indeed, the reason is that the states of the system cannot be chosen freely, but depend on the choice of inputs and the (unknown) dynamics (10). Therefore, it is in general not possible to design an experiment so that the kernel of $\bar{D}$ coincides with a given subspace. We do note, however, that the problem simplifies in the case that $N=n+m$. In this case, if $\operatorname{rank}(\bar{D})=\operatorname{rank}(\bar{F})=n+m$, then $\ker(\bar{D})=\{0\}$ and thus $\operatorname{rank}(\bar{D}\bar{F})=n+m$. This suggests that it is important to design experiments of length precisely $N=n+m$, which will be explored in Section 5. In what follows, we summarize our progress in a proposition.
Proposition 1.
Let Assumption 2 hold. If $N=n+m$, $\operatorname{rank}(\bar{F})=n+m$ and
$\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,N-1]}\\ \mu_{[0,N-1]}\end{bmatrix}\right)=n+m,$ (36)
then (17) holds.
Example 3.
(Rank of $\bar{F}$ for various filter functions). Let $N=M=n+m$. For the function $f_{\ell}$ in (27), $\bar{F}=I.$ For the function $f_{\ell}$ in (28),
$\bar{F}=\begin{bmatrix}1&0&\cdots&0\\ e^{-\rho T}&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ e^{-\rho(N-1)T}&e^{-\rho(N-2)T}&\cdots&1\end{bmatrix}$. (37)
Moreover, for the function $f_{\ell}$ in (29),
$\bar{F}=\begin{bmatrix}e^{-\rho T}&e^{-2\rho T}&\cdots&e^{-\rho NT}\\ 0&e^{-\rho T}&\cdots&e^{-\rho(N-1)T}\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&e^{-\rho T}\end{bmatrix}$. (38)
Note that $\operatorname{rank}(\bar{F})=n+m$ in all three cases.
As shown in Example 3, all principal submatrices of $\bar{F}$ derived from (27)-(29) are full rank. This observation, together with (32), leads to a rank relation between the filtered data and the sampled data.
Proposition 2.
Consider a trajectory $(u,x)\in\mathfrak{B}_{NT}$. Let the functions $g_{\ell}$, for $\ell=1,2,\dots,M$, satisfy Assumption 2 with $f_{\ell}((\ell-1)T)\neq 0$ and $f_{\ell}(kT)=0$ for all $k\geq\ell$ and $\ell\in\{1,2,\dots,\min\{N,M\}\}$. Then,
$\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,k-1]}\\ \mu_{[0,k-1]}\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}x^{\rm f}_{[1,k]}\\ u^{\rm f}_{[1,k]}\end{bmatrix}\right),$ (39)
for all $k\in\{1,2,\dots,\min\{N,M\}\}$.
Proof. Since $f_{\ell}(kT)=0$ for all $k\geq\ell$, by considering the first $k$ columns of (32), we have
$\begin{bmatrix}x^{\rm f}_{[1,k]}\\ u^{\rm f}_{[1,k]}\end{bmatrix}=\bar{C}\begin{bmatrix}\chi_{[0,k-1]}\\ \mu_{[0,k-1]}\end{bmatrix}\bar{F}_{k},$ (40)
where
$\bar{F}_{k}=\begin{bmatrix}f_{1}(0)&f_{2}(0)&\cdots&f_{k}(0)\\ 0&f_{2}(T)&\cdots&f_{k}(T)\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&f_{k}((k-1)T)\\ \end{bmatrix}$.
Since $f_{\ell}((\ell-1)T)\neq 0$ for all $\ell\in\{1,2,\dots,k\}$, $\bar{F}_{k}$ is nonsingular. By Lemma 3, $\bar{C}$ is also nonsingular. Hence, (32) implies that (39) holds for any $k\in\{1,2,\dots,\min\{N,M\}\}$. $\Box$
In the next section, we will focus on an experiment design method that imposes the rank condition (36) by appropriate choice of the inputs. This, in combination with Proposition 2, will lead to the desired rank condition on the filtered data as in (17).
5 Online experiment design
The basic idea to achieve (36) involves designing the input so that the rank of the data matrix increases progressively, that is, for all $k\in\{1,2,\dots,n+m-1\}$,
$\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,k-1]}&\chi_{k}\\ \mu_{[0,k-1]}&\mu_{k}\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,k-1]}\\ \mu_{[0,k-1]}\end{bmatrix}\right)+1.$ (41)
Condition (41) can be satisfied if the input signal is designed during system operation via the following online procedure.
Proposition 3.
Let Assumption 1 hold. For the continuous-time system (10), design the piecewise constant input signal $u$ in (19) as follows. At time $t=0$, select a nonzero $\mu_{0}$. At time $t=kT$, for $k\in\{1,2,\dots,N-1\}$,
  • if
    $\chi_{k}\notin\operatorname{im}\chi_{[0,k-1]},$ (42)
    then select $\mu_{k}$ arbitrarily;
  • if
    $\chi_{k}\in\operatorname{im}\chi_{[0,k-1]},$ (43)
    then there exist $\eta\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{n}$ with $\eta\neq 0$ such that
    $\begin{bmatrix}\xi^{\top}&\eta^{\top}\end{bmatrix}\begin{bmatrix}\chi_{[0,k-1]}\\ \mu_{[0,k-1]}\end{bmatrix}=0.$ (44)
    In this case, select $\mu_{k}$ such that
    $\xi^{\top}\chi_{k}+\eta^{\top}\mu_{k}\neq 0.$
Then, the resulting trajectory $(u,x)\in\mathfrak{B}_{(n+m)T}$ satisfies (36).
Proposition 3 follows from [van2021beyond, Theorem 1] applied to the discrete-time system (22). Its rationale is as follows. For a piecewise constant input signal $u$, the sampled data obey the dynamics in (22). Then, by Assumption 1 and Lemma 1, at time instant $kT$, $k\in\{0,1,\dots,N-1\}$, a $\mu_{k}$ can always be designed such that (41) holds.
By combining Proposition 1 and Proposition 3, we conclude that Problem 2 is solved. We summarize this in the following theorem.
Theorem 1.
Let Assumption 1 and 2 hold, and suppose that $\operatorname{rank}(\bar{F})=n+m$. Let the input signal $u$ be designed according to Proposition 3. Then, the resulting trajectory $(u,x)\in\mathfrak{B}_{\mathcal{T}}$ is such that the filtered data $(u^{f}_{[1,M]},x^{f}_{[1,M]})$ satisfy (17).
Let us now go one step further. The following result shows that the online input design ensures not only that the rank condition in (36) holds, but that the stronger condition in (25) holds.
Theorem 2.
Let Assumption 1 hold. Consider the trajectory $(u,x)\in\mathfrak{B}_{NT}$ where $u$ is a piecewise constant input signal designed as in Proposition 3. Then, (25) holds.
Proof. Let $t\in[0,T)$, and $\eta\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{n}$ be vectors such that
$\begin{bmatrix}\xi^{\top}&\eta^{\top}\end{bmatrix}\begin{bmatrix}\chi_{[0,N-1]}(t)\\ \mu_{[0,N-1]}\end{bmatrix}=0.$ (45)
By (10), for any $k\in\{1,2,\dots,N-1\}$ we get
$\chi_{k}(t)$ $=e^{At}\chi_{k}+\int_{kT}^{t+kT}e^{A(t+kT-\tau)}Bu(\tau)d\tau$ (46)
$=e^{At}\chi_{k}+\int_{0}^{t}e^{A(t-\tau)}Bd\tau\mu_{k}.$
Then, we have
$\chi_{[0,N-1]}(t)=\begin{bmatrix}e^{At}&\int_{0}^{t}e^{A(t-\tau)}Bd\tau\end{bmatrix}\begin{bmatrix}\chi_{[0,N-1]}\\ \mu_{[0,N-1]}\end{bmatrix}$, (47)
implying that
$\begin{bmatrix}\xi^{\top}&\eta^{\top}\end{bmatrix}\begin{bmatrix}e^{At}&\int_{0}^{t}e^{A(t-\tau)}Bd\tau\\ 0&I\end{bmatrix}\begin{bmatrix}\chi_{[0,N-1]}\\ \mu_{[0,N-1]}\end{bmatrix}=0.$ (48)
According to Proposition 3, (36) holds, implying that
$\xi^{\top}e^{At}=0\text{ and }\xi^{\top}\int_{0}^{t}e^{A(t-\tau)}Bd\tau+\eta^{\top}=0.$ (49)
Since $e^{At}$ is nonsingular, we have $\xi=0$, implying $\eta=0$. Hence, we conclude that (25) holds. $\Box$
Next, we compare Theorem 2 to Lemma 2, i.e., [lopez2022continuous, Lemma 1]. It is not difficult to see that Lemma 2 requires at least $n+m+nm$ samples to guarantee the full rank condition (25). The importance of Theorem 2 is to guarantee (25) with the minimum number of samples possible, namely, $n+m$. We further note that differently from Lemma 2, where the input signal can be designed entirely before data is collected from the system, the experiment design procedure in Proposition 3 constructs the input signal during system operation.
6 Numerical examples
In this section, we consider numerical examples of the rank condition relating sampled data to filtered data, and we demonstrate the sample efficiency of the proposed online experiment design.
Consider the aircraft longitudinal dynamics [moustakis2018adaptive] with
$A=$ $\begin{bmatrix}-0.0190&0.0825&-0.1005&-0.3206\\ -0.2154&-2.7859&1.2031&-0.0271\\ 3.2527&-30.7871&-3.5418&0\\ 0&0&1&0\end{bmatrix}$,
$B=$ $\begin{bmatrix}0.0065&0.0534\\ -0.6103&0.0020\\ -74.6355&0.5431\\ 0&0\end{bmatrix}$,
where the four states are the velocity along the $x$- and $z$-body-axis, the angular velocity around the $y$-body-axis, and the pitch angle, while the two inputs are the elevator deflection and throttle. Note that $n+m=6$. Let $T=0.1$. Because the eigenvalues of $A$ are $\{-3.1548\pm 6.0892i,-0.0185\pm 0.3263i\}$, Assumption 1 holds. In general, although $A$ is unknown, Assumption 1 fails to hold only on a set of measure zero. Let $x(0)=\begin{bmatrix}2&-1&1&0.5\end{bmatrix}^{\top}$. By designing the input as in Proposition 3, we collect the input-state samples
$\chi_{[0,6]}=$
$\begin{bmatrix}2&1.9877&1.9308&1.8672&1.7965&1.7030&1.6170\\ -1&-0.9492&-0.4078&-0.0961&0.2965&0.6994&0.7164\\ 1&-2.5648&6.9073&-0.5012&6.2982&3.5254&0.9785\\ 0.5&0.4124&0.6720&0.9783&1.2988&1.7922&2.0105\end{bmatrix}$, (50)
$\mu_{[0,5]}=\begin{bmatrix}1&-1&1&-1&0&0\\ 1&-1&0&0&1&-1\end{bmatrix}$. (51)
As specific examples of filter functions, let us consider the compactly supported test function in (2) and the low-pass filter function in (5) with $\rho=1$.
For the compactly supported test function, we obtain from (13), (14) and (15) the following data matrices
$x^{\rm f}_{[1,6]}=$
$\!\!\begin{bmatrix}0.6632&0.6566&0.6306&0.6133&0.5827&0.5527\\ -0.3090&-0.2614&-0.0489&0.0081&0.1820&0.2471\\ -0.3009&0.9059&1.0074&1.0977&1.6470&0.7208\\ 0.1649&0.1468&0.3017&0.3552&0.5252&0.6429\end{bmatrix}\!$, (52)
$u^{\rm f}_{[1,6]}=$
$\!\!\begin{bmatrix}0.3333&-0.3333&0.3333&-0.3333&0&0\\ 0.3333&-0.3333&0&0&0.3333&-0.3333\end{bmatrix}\!$, (53)
$x^{\rm df}_{[1,6]}=$
$\!\!\begin{bmatrix}-0.0407&-0.1921&-0.2118&-0.2373&-0.3122&-0.2865\\ 0.1488&1.8755&1.0010&1.3597&1.3355&0.0418\\ -11.9604&31.6734&-24.8886&22.7354&-9.3601&-8.5424\\ -0.3009&0.9058&1.0074&1.0977&1.6470&0.7208\end{bmatrix}\!$. (54)
For the low-pass filter function, we obtain from (13), (14) and (15) the following data matrices
$x^{\rm f}_{[1,6]}=$
$\!\!\begin{bmatrix}0.1894&0.3586&0.5046&0.6314&0.7377&0.8252\\ -0.0892&-0.1527&-0.1541&-0.1352&-0.0711&0.0055\\ -0.0862&0.1766&0.4453&0.7133&1.1128&1.2126\\ 0.0462&0.0861&0.1625&0.2503&0.3761&0.5235\end{bmatrix}$, (55)
$u^{\rm f}_{[1,6]}=$
$\!\!\begin{bmatrix}0.0952&-0.0091&0.0870&-0.0165&-0.0149&-0.0135\\ 0.0952&-0.0091&-0.0082&-0.0074&0.0885&-0.0151\end{bmatrix}$, (56)
$x^{\rm df}_{[1,6]}=$
$\!\!\begin{bmatrix}-0.0114&-0.0653&-0.1190&-0.1756&-0.2477&-0.3058\\ 0.0449&0.5636&0.7988&1.1021&1.3770&1.2597\\ -3.3835&5.9120&-1.6873&4.9145&1.8061&-0.7829\\ -0.0862&0.1765&0.4453&0.7133&1.1128&1.2126\end{bmatrix}$. (57)
For both filter functions, we verify that
$\operatorname{rank}\left(\begin{bmatrix}x^{\rm f}_{[1,6]}\\ u^{\rm f}_{[1,6]}\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,5]}\\ \mu_{[0,5]}\end{bmatrix}\right)=6.$ (58)
This is in line with the conclusion of Proposition 1. Finally, we reconstruct the system matrices as in (18)
$\begin{bmatrix}\hat{A}&\hat{B}\end{bmatrix}=x^{\rm df}_{[1,6]}\begin{bmatrix}x^{\rm f}_{[1,6]}\\ u^{\rm f}_{[1,6]}\end{bmatrix}^{-1}.$
The Frobenius norm of the identification error is $\lVert[A\ B]-[\hat{A}\ \hat{B}]\rVert_{F}=6.2340\times 10^{-7}$ for the compactly supported test function, and $1.4599\times 10^{-6}$ for the low-pass filter function. We mention that similar conclusions can be drawn for the filter functions in (3) and (4), which are not shown for the sake of brevity.
7 Conclusion
This paper has investigated the problem of experiment design for continuous-time systems. To avoid reliance on time derivatives of measured trajectories, which is a key challenge in data-driven methods for continuous-time systems, we have proposed a generalized filtering framework to collect filtered data. We have presented conditions to ensure that these filtered data are informative for system identification. We then have developed an online experiment design method that guarantees informativity with the minimum number of samples. Several examples of filter functions have demonstrated the generality of the proposed filtering framework. Potential future work includes applying this experiment design methods in specific data-driven control settings such as data-driven stabilization or model reference control.